y=10sinpie[0.01x-2.00t]find maximum speed of tranverse wave

  1. y=10sinpie[0.01x-2.00t]find maximum speed of tranverse wave

Grade:12th pass

1 Answers

Arun
6 years ago
Given the partial derivative WRT t found earlier we have Partial dy/dt= -20pi*cos(pi(.01x - 2.00t)) Now I understand that the maximum will occur when the trig portion is equal to 1. That leaves me with only the -20pi
Hence Maximum speed = 20* pi = 20 * 3.14 = 62.8 m / sec
Regards
Arun (askIITians forum expert)